### v2

# Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models

Abstract
Recent Large Reasoning Models (LRMs) excel at complex reasoning tasks but often suffer from overthinking, generating overly long and redundant reasoning trajectories. To explore its essence, our empirical analysis reveals that LRMs are primarily limited to recognizing task properties (i.e., difficulty levels) like humans before solving the problem, leading to a one-size-fits-all reasoning process. Inspired by this, a pressing and natural question emerges: Can we explicitly bootstrap such ability to alleviate overthinking in LRMs? In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty cognition and redundancy cognition of LRMs. Specifically, we first inject difficulty hypnosis into output prefixes to guide the model toward adaptive reasoning depth, trained on a hybrid dataset mixing short and long reasoning paths. Then, we incorporate redundancy hypnosis, which supervises the intermediate reasoning steps to identify and eliminate unnecessary reasoning patterns. Experiments on 7B/14B/32B models demonstrate that TH2T significantly reduces inference costs by over 70% on easy tasks and 40% on hard tasks while maintaining performance stability. The resulting outputs exhibit clear signs of difficulty-aware capabilities and reduced redundancy (e.g., reflection and looping).
Introduction
The emergence of reasoning-focused Large Language Models, also known as Large Reasoning Models (LRMs), has led to substantial improvements in problem-solving capabilities (Sun et al. 2024b; Yu et al. 2024; Team et al. 2025), particularly through the use of elaborate and sequential reasoning enabled by test-time scaling (Snell et al. 2025). In addition to OpenAI-O1 (OpenAI et al. 2024a), other models such as the open-source Deepseek-R1 (Guo et al. 2025) and QwQ (Team 2025) leverage long Chain-of-Thought (Wei et al. 2022; Wang et al. 2023) and adopt reinforcement learning (RL) to approach complex tasks step-by-step, enabling a human-like thinking pattern with reflection, backtracking, and self-validation (Guo et al. 2025; Xie et al. 2025; Liu et al. 2025). This reasoning paradigm demonstrably enhances the models’ problem-solving capabilities and yields promising results. However, lengthy responses introduce the problem of “overthinking” (Team et al. 2025; Chen et al. 2025), wherein model performance initially improves with extended reasoning chains but deteriorates beyond a certain length. For instance, even for trivial queries such as “What is 1 plus 1?”, current models often misjudge the difficulty and produce unnecessarily long outputs—sometimes spanning hundreds of tokens. This tendency toward redundant reasoning not only increases computational cost and latency but also degrades user experience. These observations prompt a critical question: do LRMs possess the capacity to recognize and adapt to varying levels of problem difficulty autonomously?
To answer this question, we start from the perspective of cognitive science, specifically the dual-process theory of human reasoning. This theory posits that human cognition operates via two distinct but complementary systems: System 1, which is fast, concise, and intuitive, and System 2, which is slow, deliberate, and analytical (Kahneman 2011; Evans and Stanovich 2013). Before engaging in problem-solving, humans often perform an initial assessment of a task’s difficulty, allowing them to allocate cognitive resources appropriately between these two systems (Efklides 2011; Flavell 1979). For straightforward tasks, reliance on System 1 suffices, whereas complex problems necessitate the engagement of System 2. This metacognitive ability to evaluate task difficulty and adapt reasoning strategies accordingly, however, is lacking in current reasoning models. As shown in Fig. 2, DeepSeek-R1-Distill-Qwen2.5-7B frequently conflates easy datasets (e.g., GSM8K (Cobbe et al. 2021)) and hard datasets (e.g., AIME2024), treating them with a uniform level of reasoning complexity. These limitations highlight the need for LRMs to develop autonomous difficulty cognition—to “think how to think”.
Current research primarily emphasizes unified CoT compression based on length, yet length is not an accurate proxy for difficulty. In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs difficulty cognition and redundancy cognition. In contrast to prompt-based input guidance for instruction adherence, our approach emphasizes output-level interventions, directly intervening inside the reasoning trajectory to establish a self-hypnosis mechanism. Specifically, in the first stage, to encourage the model to develop native difficulty cognition autonomously, we design difficulty-awareness hypnosis and inject it into response prefixes to establish cognitive hints. Combined with a heterogeneous dataset, the trained model enhances its sensitivity to task difficulty, thereby reducing overthinking in simpler tasks through structured cognitive self-guidance.
In the second stage, we conduct fine-grained in-progress redundancy management, enhancing LRM’s redundancy cognition by targeting redundant yet logically complete inference structures (e.g., reflections or double-checks) and terminal reasoning loops commonly observed in complex task generations. The identified structures are injected with redundancy hypnosis while truncating subsequent superfluous content. Overall, our contributions can be outlined as follows:
We observe that current LRMs fall short of perceiving task complexity. To mitigate overthinking, we propose enhancing LRMs with the ability to cognize and adapt to varying levels of problem difficulty autonomously.
To align with human System 1 and System 2 thinking patterns, we propose a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty and redundancy cognition with a constructed heterogeneous dataset with inserted self-hypnosis.
Experiments demonstrate over 70% length reduction on easy tasks and 40% on hard tasks with maintained performance stability, exhibiting clear difficulty-aware capabilities and reduced redundancy.
Literature Review
The advent of LRM (OpenAI et al. 2024a; Guo et al. 2025; Team 2025) enables breakthroughs in problem solving (Sun et al. 2024b; Yu et al. 2024) with Chain-of-Thought (Wei et al. 2022; Wang et al. 2023) and test-time scaling (Snell et al. 2025). Nevertheless, this introduces the challenge of overthinking (Chen et al. 2025; Team et al. 2025). Works of efficient reasoning (Sui et al. 2025) as summarized:
Model-based methods.
Reinforcement Learning motivates some existing efficient reasoning works (Arora and Zanette 2025; Team et al. 2025; Aggarwal and Welleck 2025) to incorporate length penalties. O1-pruner (Luo et al. 2025b) introduces the Length-Harmonizing Reward, integrated with a PPO-style loss. DAST (Shen et al. 2025b) constructs a length-preference dataset and employs SimPO (Meng and Xia 2024). AdaR1 (Luo et al. 2025a) constructs a hybrid reasoning model by merging long and short CoT models and is optimized with DPO (Rafailov et al. 2023). Apart from RL, which is vulnerable and expensive, fine-tuning is another effective post-training strategy (Yang et al. 2025b). CoT-Value (Ma et al. 2025b) regulates reasoning length by a controllable direction in parameter space, enabling dynamic length control. C3oT (Kang et al. 2025) employs GPT-4 as a compressor while retaining key information. TokenSkip (Xia et al. 2025) prunes less important tokens and fine-tune the model to skip the redundant tokens.
Output-based methods.
Implicit-based methods (Deng, Choi, and Shieber 2024; Hao et al. 2024; Saunshi et al. 2025; Cheng and Durme 2024; Shen et al. 2025a) propose to internalize reasoning steps within the latent representation, transitioning reasoning from language space to hidden space. (Saunshi et al. 2025) simulates multiple steps of CoT through a looped transformer layer, increasing the depth of reasoning without additional parameters. However, these implicit techniques introduce performance decline and affect interpretability. There are also dynamic criteria. For example, (Sun et al. 2024a; Xie et al. 2023; Liao et al. 2025) optimize Best-of-N (BoN) decoding with speculation and rejection. (Li et al. 2024; Yang et al. 2025a; Wang et al. 2025) explores possible early exit point. NoThinking (Ma et al. 2025a) prompts reasoning models to directly output final answers without thinking, but heavily relies on sampling, which adversely increases overall overhead.
Prompt-based methods.
Some works (Aytes, Baek, and Hwang 2025; Ding et al. 2024; Ning et al. 2024; Han et al. 2025; Xu et al. 2025; Lee, Che, and Peng 2025) instruct LRMs by various prompts to respond concisely. Specifically, (Renze and Guven 2024) explicitly prompts LLMs to perform concise step-by-step reasoning. SoT (Ning et al. 2024) combines reasoning paradigms with linguistic constraints to minimize token usage. Token-Budget (Han et al. 2025) establishes a token budget estimated by LLM itself, and incorporates this estimation as a token number constraint. (Ge et al. 2025) regulates thinking token distribution by few-shot in-context learning. Nevertheless, prompting relies on precise instruction following and robust task cognition, which will be proven unreliable later.
Overall, most existing works treat efficient reasoning as a unified length compression paradigm. There is a necessity that has been widely overlooked that LRM may need the capacity to autonomously recognize and adapt to varying levels of problem difficulty, which aligns with System 1 and System 2 thinking patterns of human cognitive theory.
Preliminary and Observation
Background and Notations
LLMs.
Given a prompted question , conventional LLM generates a response by iteratively predicting each response token conditioned on and generated tokens . This is termed as System 1 (Li et al. 2025b) thinking:
where [Prompt] and [Answer] are question and response , which is often short and concise.
LRMs.
Whereas in reasoning-focused LRMs, the generation process is split into two stages, e.g., thinking and conclusion, which is referred to as System 2 (Li et al. 2025b):
where and are two delimiters denoting the start and end of reasoning, [Thoughts] and [Conclusion] are detailed reasoning process and concise answer, respectively. Formally, the reasoning part [Thoughts] reveals a trajectory of problem solving, known as CoT (Wei et al. 2022; Wang et al. 2023). At the micro level, some temporally consecutive tokens in [Thoughts] can be aggregated into short chunk based on semantics, logic and grammar:
Similarly, each reasoning chunk in this paradigm is autoregressively generated by conditioning on question and previous reasoning thoughts: . This explicit decomposition not only improves interpretability through transparent exposure of reasoning steps but also enhances the model’s capability for complex tasks. Specifically, some chunks are well structured by prefix keywords such as “Wait”, “However”, “Alternatively” (Guo et al. 2025), marking the subsequent chunks will undergo logical backtracking, self-verification, or exploration. Works (Li et al. 2025a) have shown that reflection structures are crucial to LRMs. Under this setting, the “overthinking” phenomenon is that many seem delicate yet contribute nothing to reasoning logic, resulting in prominent overhead.
Observation: Reasoning Process Intervention
To address the overthinking issue, a straightforward approach is difficulty reminders—encouraging LRMs to self-regulate output length based on task difficulty judgment through prompts like “Please think quickly if you find it easy”. However, this conventional interventional technique necessitates two prerequisites: accurate instruction following and robust task difficulty cognition. While prior studies (Wu et al. 2025; Han et al. 2025) have highlighted the inherent challenges of the former, we have already shown that current LRMs struggle to recognize varying levels of problem difficulty in Fig. 2.
Instead of using a difficulty reminder in the prompt, we claim that a promising solution is transferring the intervention from the prompt to the inner reasoning process. We conduct a pre-experiment in Fig. 3, pretending the model has already correctly recognized the difficulty level of samples. Specifically, for a simple problem, we change the model input from “User: [Prompt]+[Question]. Assistant:” to “User: [Prompt]+[Question]. Assistant: This is a simple question, let’s think quickly.”. While for hard problems, the corresponding predefined response prefix is “It seems difficult, let’s think thoroughly.”. Fig. 3 shows that predefined difficulty awareness indeed helps to reduce response length, especially matched awareness (“simple, quickly” for easy and “difficult, thoroughly” for hard is matched, the reverse is unmatched).
We define them as self-hypnosis that acts as a trigger in the output, guiding the subsequent inference trajectory. In fact, the reflection structures in LRMs (Guo et al. 2025; Li et al. 2025a) are essentially similar triggers aiming to intervene in subsequent think chunks. Compared to explicit prompts, we argue that these built-in triggers offer three key advantages: (1) native flexibility, (2) robustness, and (3) superior instruction-following capabilities. In the next section, we introduce how we guide LRMs to generate native interventional triggers of difficulty and redundancy cognition, autonomously.
Methodology
In this section, we describe our two-stage fine-tuning strategy Think-How-to-Think (TH2T) for mitigating overthinking, as shown in Fig. 4.
Stage 1: Difficulty Cognition Injection
We aim to endow LRM with the capability to assess the difficulty of a given problem. We define two datasets: for easy problems, and for difficult ones. Two models from the same series were selected to ensure consistency in response: for basic LLM with prompted Short-CoT capability (e.g., Qwen2.5 series) and for advanced LRM (e.g., DeepSeek-R1-Distill-Qwen2.5 series). For each problem in , the responses are recorded as:
where means the response of on is correct. can replicate the majority correct responses of on , which means: , wherein significant additional overhead yields limited performance gain. This motivates us to preserve ’s potential for efficiency on with satisfactory performance. Our desired thinking pattern of is responsed concisely on and elaborately on . To align with this objective, we construct a draft dataset grounded in difficulty differentiation:
where , similar to Eq. (5). aims to maintain the performance gain of on .
To enable to perceive problem difficulty level autonomously, we introduce the trials of Fig. 3. We inject specific difficulty-aware cognitive Hypnosis into the prefix of answer within , which, as demonstrated before, functions as interventional triggers guiding the reasoning trajectory:
If , the difficulty hypnosis H is defined as follows: “hypnosis This is a simple question, let’s think quickly./hypnosis”. If , then “hypnosisIt seems difficult, let’s think thoroughly./hypnosis”. The markers “hypnosis” and “/hypnosis” are defined as indicators for inserted difficulty hypnosis. The injected hypnosis is recognized as an embedded difficulty label to train to generate native difficulty cognition on the refined dataset , stimulating a concise thinking pattern on easy tasks. The execution of stage 1 is outlined at the upper section of the algorithmic framework Fig. (4).
Stage 2: Redundancy Cognition Injection
We extend difficulty-hypnosis of Stage 1 to redundancy-hypnosis, guiding to identify redundant structures within the reasoning process on hard tasks.
[…]
where is the difficulty-hypnosis introduced in Stage 1, are valid reasoning chunks to be kept, is the unnecessary reflection structure inside the reasoning process without effective contribution to the reasoning process. These chunks often carry out backtracking, self-verification, or exploration, but unfortunately, are nothing more than superfluous overhead. is a similar reflection structure; nonetheless, it primarily manifests in the final phase of the reasoning process, continuously iterating until attains the maximum output length.
We define a Determinator to judge whether some is logically equivalent to or (i.e., redundant or looping). If yes, we inject redundancy hypnosis and truncate subsequent content. In case of , original is refined as:
wherein the redundancy hypnosis is defined as follows: “hypnosis Everything seems ok, let’s move on./hypnosis”. We use GPT-4 (OpenAI et al. 2024b) to implement the determinator.
In case of , on one hand, the training set lacks a sufficient number of loop structures; on the other hand, excessively deferred looping hinders the generation of shorter-length samples. Thus, we select intermediate chunk with words like “Wait”, “However”, “Alternatively” and iterate it as a simulation of looping, followed by hypnosis injection and truncation:
wherein the looping hypnosis is defined as follows: “hypnosis Oh, I’m stuck in a loop. Time to break out. /hypnosis”. The choice of iterate number can be random, and we also try a probe-like approach (Zhang et al. 2025). We use DeepSeek-R1 to identify chunks before which the prior content can already induce the correct answer, referred to as completeness requirements. The last-layer hidden states of the selected chunk tail is extracted to train a simple two-layer MLP as Determinator. Experiments indicate it yields similar results to the random approach, so we prefer the latter.
Notably, we adopt truncation instead of mere substitution, as truncation allows for the construction of shorter samples. Meanwhile, generation continuity is ensured by remaining untruncated samples in the dataset, since we only apply Stage 2 to part of the samples in . Stage 2 execution is outlined in Fig. 4.
Experiments
Experiment Setup
Models.
Dataset Construction.
For dataset construction, we mainly adopt mathematical reasoning tasks, which are friendly for difficulty quantization. In stage 1, the GSM8K and MATH datasets are combined after filtering out errors and injecting difficulty hypnosis. In stage 2, redundancy and looping hypnosis are injected into difficult samples. The composition ratio of this heterogeneous dataset is 2:1:1.
Baselines.
We compare TH2T with two prompt-based methods, i.e., D-Prompt and NoThinking (Ma et al. 2025a), and two fine-tuning methods, i.e., Tokenskip (Xia et al. 2025), CoT-Valve (Ma et al. 2025b). D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty, through difficulty reminder in the prompt, as discussed in Preliminary Part. The remaining methods are from recent publications.
| Methods | GSM8K | MATH-500 | ||||||||||
| Acc | Gain | Len | Reduc | Latency | SpdUp | Acc | Gain | Len | Reduc | Latency | SpdUp | |
| DeepSeek-R1-Distill-Qwen-7B | ||||||||||||
| Original | 91.51 | 1057 | 20.29 | 85.2 | 2857 | 87.69 | ||||||
| [1pt/2pt] D-Prompt | 90.59 | 0.92 | 906 | 14.3% | 16.38 | 1.24 | 86.0 | 0.8 | 2726 | 4.6% | 81.54 | 1.08 |
| NoThinking | 88.47 | 3.04 | 243 | 76.9% | 3.34 | 6.07 | 79.4 | 5.8 | 699 | 75.5% | 11.81 | 6.43 |
| [1pt/2pt] TokenSkip | 88.01 | 3.50 | 638 | 39.6% | 10.64 | 1.91 | 80.2 | 5.0 | 2109 | 26.2% | 53.94 | 1.62 |
| CoT-Valve | 87.90 | 3.61 | 235 | 77.8% | 3.31 | 6.13 | 78.6 | 6.6 | 747 | 73.9% | 12.31 | 7.12 |
| (w/o D.H.) | 90.83 | 0.68 | 503 | 47.4% | 7.80 | 2.60 | 85.2 | 0.0 | 1906 | 33.2% | 46.68 | 1.88 |
| (w/o R.H.) | 91.96 | 0.48 | 397 | 62.4% | 5.99 | 3.39 | 86.0 | 0.8 | 2265 | 20.7% | 60.22 | 1.46 |
| TH2T | 92.87 | 1.36 | 275 | 74.0% | 3.88 | 5.23 | 86.8 | 1.6 | 1772 | 38.0% | 41.65 | 2.11 |
| DeepSeek-R1-Distill-Qwen-14B | ||||||||||||
| Original | 93.63 | 584 | 9.38 | 86.6 | 2306 | 62.65 | ||||||
| [1pt/2pt] D-Prompt | 93.91 | 0.28 | 606 | 3.9% | 9.95 | 0.94 | 86.6 | 0.0 | 2297 | 0.3% | 62.60 | 1.01 |
| NoThinking | 88.41 | 3.48 | 206 | 64.9% | 3.06 | 3.07 | 76.0 | 10.6 | 612 | 73.5% | 10.05 | 6.23 |
| [1pt/2pt] TokenSkip | 89.30 | 4.33 | 314 | 46.2% | 4.79 | 1.96 | 73.2 | 13.4 | 1356 | 41.2% | 28.11 | 2.23 |
| CoT-Valve | 92.70 | 0.90 | 526 | 9.9% | 8.13 | 1.15 | 84.2 | 2.4 | 2111 | 8.4% | 53.95 | 1.16 |
| (w/o D.H.) | 93.69 | 0.06 | 338 | 42.2% | 4.82 | 1.95 | 86.2 | 0.4 | 1549 | 32.8% | 33.93 | 1.85 |
| (w/o R.H.) | 93.47 | 0.16 | 299 | 51.2% | 4.34 | 2.16 | 85.8 | 0.8 | 1706 | 26.0% | 39.83 | 1.57 |
| TH2T | 94.08 | 0.45 | 246 | 58.1% | 3.38 | 2.78 | 86.4 | 0.2 | 1343 | 41.8% | 27.84 | 2.25 |
| DeepSeek-R1-Distill-Qwen-32B | ||||||||||||
| Original | 95.60 | 717 | 12.14 | 87.2 | 2357 | 63.02 | ||||||
| [1pt/2pt] D-Prompt | 95.60 | 0.00 | 663 | 7.6% | 11.12 | 1.09 | 86.8 | 0.4 | 2198 | 6.8% | 57.11 | 1.10 |
| NoThinking | 93.86 | 1.74 | 228 | 68.2% | 3.25 | 3.74 | 80.2 | 7.0 | 659 | 72.0% | 10.98 | 5.74 |
| [1pt/2pt] TokenSkip | 96.54 | 0.94 | 423 | 41.0% | 6.28 | 1.93 | 79.8 | 7.4 | 1567 | 33.5% | 34.26 | 1.84 |
| CoT-Valve | 96.10 | 0.50 | 781 | 8.9% | 12.56 | 0.97 | 85.3 | 1.9 | 2263 | 3.9% | 60.21 | 1.05 |
| (w/o D.H.) | 96.08 | 0.48 | 361 | 49.7% | 5.47 | 2.22 | 87.4 | 0.2 | 1863 | 20.9% | 45.18 | 1.39 |
| (w/o R.H.) | 95.75 | 0.15 | 323 | 55.0% | 4.84 | 2.51 | 86.8 | 0.4 | 2114 | 10.3% | 53.99 | 1.17 |
| TH2T | 97.04 | 1.44 | 263 | 63.3% | 3.51 | 3.46 | 88.0 | 0.8 | 1733 | 26.5% | 40.13 | 1.57 |
| Bench | Difficulty Level | Cognition | |
| Original | TH2T | ||
| GSM8K | all easy | 14.9% | 98.3% |
| MATH | lv.1-2 (easy) | 9.8% | 93.4% |
| lv.4-5 (hard) | 12.0% | 91.5% | |
| AIME2024 | all hard | 16.7% | 96.7% |
| OmniMath | lv.1-2 (easy) | 20.0% | 96.7% |
| lv.4-5 (hard) | 13.3% | 93.3% | |
Evaluations.
We evaluate on 4 mathematical benchmarks (GSM8K, MATH, AIME2024 and Omni-MATH) and 1 QA benchmark GPQA (Rein et al. 2024). Metrics include accuracy, average number of output tokens, and inference latency per sample. We also report the success rate of the difficulty awareness, redundancy, and looping optimization.
The detailed experiment settings are in the appendix.
Experimental Results
Main Results.
The performance of TH2T on GSM8K and MATH is illustrated in Tab. 1. Key observations include: 1) TH2T exhibits minimal accuracy degradation; more precisely, most settings witness slight improvement. 2) TH2T achieves the best reduction in reasoning length among methods with stable accuracy—up to 74% on GSM8K and 38% on MATH with the 7B model, translating to over 5 and 2 latency speedups, respectively. 3) Greater gains on easier tasks, validating TH2T’s ability to adapt reasoning effort based on task difficulty. 4) These observations are consistent across 7B, 14B and 32B models. As shown in Fig. 1, other methods (e.g., NoThinking and CoT-Value) achieve comparable compression ratios in token length but at the cost of severe accuracy degradation (often 10 points), underscoring TH2T’s robustness in balancing efficiency and reliability.
Difficulty Cognition.
Does TH2T present better difficulty cognition?. As shown in Tab. 2, the entirety of GSM8K, MATH lv.1-2, and Omni-Math lv.1-2 are set as “Easy” while the remaining are set as “Hard”. The original R1-Distilled-7B model is prompted to predict its own understanding of problem difficulty and choose between “Easy”, “Medium”, and “Hard”. TH2T can response with native self-hypnosis in the prefix like “This is a simple question” or “It seems difficult”, which indicates the cognition of task difficulty as “Easy” or “Hard”. As shown in Tab. 2 and Fig. 2, the original LRM demonstrates ambiguous recognition of task difficulty, uniformly categorizing tasks of varying complexity as medium level. Whereas, our approach exhibits a clear distinction with high precision over 90+%, validating the effectiveness of the first stage of TH2T.
Redundancy Cognition.
Does TH2T present better redundancy cognition? Tab. 3 shows the average number of reflective and looping structures. Comparison is made between R1-Distilled-7B and Ours-7B. We observe that: 1) TH2T effectively suppresses the number of reflection chunks by 10 on easy tasks and 3 on hard tasks, suggesting that extensive reflection is unnecessary for strong reasoning. 2) The looping problem in TH2T undergoes substantial restraint. Notably, 59.5% of incorrect answers by the original model on MATH exhibit tail recursion, which is mitigated to 21.4% in ours. As shown in Fig. 6(b), we observe zero loop structure in GSM8K. This distinct pattern of short-length concentration underscores the efficacy of Stage-2 of TH2T, presenting a more concise reasoning structure.
| Bench | Split | Reflective | Looping | ||
| Original | TH2T | Original | TH2T | ||
| GSM8K | Overall | 10.3 | 1.2 | 2.0% | 0.0% |
| –correct | 4.8 | 0.0 | 0.4% | 0.0% | |
| –incorrect | 69.4 | 10.8 | 22.3% | 0.0% | |
| MATH | Overall | 27.0 | 12.3 | 10.4% | 4.0% |
| –correct | 14.9 | 6.8 | 1.9% | 1.0% | |
| –incorrect | 96.5 | 39.1 | 59.5% | 21.4% | |
Generalization Ability.
To validate the generalization on OOD tasks, we evaluate on 2 more mathematical benchmarks and 1 QA benchmark. As shown in Fig. 5, accuracy maintenance and token length reduction are consistent.
Results of Llama series model are in the appendix.
Ablation Study
Ablation on Stages.
We explore two more settings: 1) without Difficulty Hypnosis in Stage-1, and 2) without Redundancy Hypnosis in Stage-2, which are annotated as w/o D.H. and w/o R.H. in Tab. 1. From the perspective of response reduction, we observe that D.H. introduces more gain on easy task GSM8K than on hard task MATH (e.g., 26.6% vs 4.8%, on 7B model), while R.H. exerts a greater impact on MATH (11.6% vs 17.3%). As we illustrate during method formulation, we design Difficulty Cognition Injection Stage to build native perception on easy tasks and Redundancy Cognition Injection Stage to mitigate non-essential structures that frequently occur in hard tasks. These results demonstrate the rationality of our two-stage strategy.
Ablation on Starting Model.
Instead of training on LRMs, we also try directly starting from a Short-CoT model with the same data in Tab. 4, which reveals a notable failure in performance and response length. This motivates us to include an underlying Stage-0 in Fig. 4 to start from an LRM, instead of teaching a Short-CoT model from scratch.
| Method | GSM8K | MATH-500 | ||
| Accuracy | Length | Accuracy | Length | |
| Qwen2.5-7B | 86.20 | 280 | 62.80 | 629 |
| [1pt/2pt] TH2T | 72.48 | 238 | 66.20 | 2333 |
| R1-Distill-7B | 91.51 | 1057 | 85.2 | 2857 |
| [1pt/2pt] TH2T | 92.87 | 275 | 86.8 | 1772 |
Ablation on Training Steps.
We examine the dynamics of accuracy and average token length during the training process (figure in appendix), and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
Conclusion
This paper presents TH2T, a novel two-stage fine-tuning strategy for large reasoning models (LRMs) that effectively addresses the challenge of overthinking. Our empirical analysis reveals that current LRMs are primarily limited to recognizing task difficulty, exhibiting a necessity for aligning with human System 1 and System 2 thinking patterns. Motivated by this, TH2T progressively enables native difficulty and redundancy cognition based on a heterogeneous dataset and proposed self-hypnosis mechanism. To the best of our knowledge, we are the first to inject difficulty/redundancy/looping hypnosises as built-in triggers for reasoning trajectory intervention. Experiments demonstrate that TH2T significantly reduces inference length with maintained performance stability, exhibiting clear difficulty and redundancy sensitivity. Future research can explore finer-grained reasoning trajectory intervention.
Supplementary Materials
More Experimental Details
Dataset Construction
The datasets used in this paper incorporate difficulty level labeling as a key consideration, so we mainly use mathematical reasoning tasks. Our TH2T is a two-stage strategy, with tailored dataset construction for each phase, followed by sequential fine-tuning. 1) During the first stage, we use the GSM8K and MATH training datasets to obtain and , respectively. The former is generally considered simple, whereas the latter is regarded as difficult and is explicitly categorized into difficulty levels ranging from 1 to 5, and we pick samples between lv.3 and lv.5. After filtering out errors, and are balanced at a 1:1 ratio. After injecting difficulty hypnosis , we construct of 6.4K samples. 2) In Stage-2, are injected with redundancy hypnosis and looping hypnosis , forming a final training dataset. We observe that the meticulous composition ratio of the constructed dataset does not significantly affect the final performance. Hence, we combine and in a ratio of 1:1. Meanwhile, the redundancy hypnosis and looping hypnosis are also injected into in a ratio of 1:1. The overall composition is set as 2:1:1.
Baselines
In our main experiments, we compare TH2T with two prompt-based methods and two fine-tuning methods.
D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty based on Difficulty Reminder in Prompt, as discussed in our Preliminary Part.
NoThinking bypasses the thinking process via prompting for all difficulty-level questions.
Tokenskip constructs a dataset where less important tokens are pruned and fine-tuned to enable the model to skip the redundant tokens in inference, regardless of task difficulty.
CoT-Valve enables dynamic control of CoT length by identifying and leveraging a controllable direction in parameter space, reducing inference cost by compressing reasoning paths based on task difficulty.
We replicate the first two baselines and report the relative performance improvement of the latter two from the raw papers due to variability of the model and implementation methods.
Benchmarks
We evaluate on 4 mathematical benchmarks and 1 QA benchmark. The main results are on the test set of in-distribution GSM8K and MATH. Meanwhile, AIME2024 and Omni-MATH are employed as out-of-distribution test sets. For Omni-MATH, we select 30 samples from difficulty ranges 1-2 and 4-5 as easy/hard instances, while AIME2024 is considered hard. To explore generalization ability in other domains, we include the GPQA benchmark to test performance on QA tasks.
Metrics
Main results are assessed on three metrics: accuracy, average number of output tokens, and inference latency per sample. To substantiate the research objectives, we incorporate the success rate of the difficulty awareness. We also report the redundancy cognition optimization, which includes the reduction of redundant structures (e.g., “Wait”, “Alternatively”, or “Check”), as well as the reflection looping ratio.
Implementation Details
We adopt LoRA, a widespread LLM fine-tuning approach, to train our models with a rank , scaling parameter , and learning rate . All experiments are conducted on 2NVIDIA H800 GPU (80GB). For the decoding strategy, we employ greedy decoding, and the maximum generation length equals 8192 (except 16,384 for AIME2024). The random seed is fixed to 42.
| Methods | GSM8K | MATH-500 | AIME2024 | OmniMath | GPQA | Average | ||||||
| Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | |
| DeepSeek-R1-Distill-Llama-8B | ||||||||||||
| Original | 75.36 | 260 | 61.8 | 701 | 3.3 | 1378 | 26.67 | 844 | 34.34 | 704 | 40.29 | 777 |
| D-Prompt | 75.14 | 261 | 62.2 | 596 | 6.6 | 1100 | 28.33 | 808 | 34.34 | 669 | 41.32 | 687 |
| NoThinking | 75.24 | 257 | 62.2 | 580 | 6.6 | 1019 | 28.33 | 714 | 32.32 | 619 | 40.94 | 638 |
| TH2T | 75.58 | 245 | 61.6 | 498 | 13.3 | 861 | 40.00 | 561 | 40.91 | 526 | 46.28 | 538 |
More Experimental Results
Results on Llama Model
We provide detailed experimental results of the R1-Distilled-Llama-8B model evaluated on 4 mathematical benchmarks (i.e., GSM8K, MATH, AIME2024, and OmniMath) and 1 QA benchmark (i.e., GPQA). We report accuracy and average token length for comparison in Tab. 5.
Acc and Length on Detailed Difficulty Levels
We report a breakdown of the accuracy and average token length distribution across the complete spectrum of difficulty levels on 4 tasks. The comparison is made between original R1-Ditstilled-Qwen–7B and TH2T-7B. Each question in the MATH and OmniMath datasets is labeled with a difficulty level, ranging from 1 to 5 and from 1 to 8, respectively. The higher the numerical level, the more difficult the question. Meanwhile, all questions in GSM8K and AIME2024 are considered easy and very difficult, respectively. Results are illustrated in Fig.7, and we obtain the following observations: 1) Both models generate progressively longer and more elaborate responses as the difficulty level increases. 2) across all difficulty levels in all tasks, TH2T exhibits significant length reduction, as well as stable performance maintenance. 3) The lower the difficulty level, the more significant the reduction in length.
Dynamics of Accuracy and Token Length
We report the dynamics of accuracy and average token length during the training steps in Fig. 8, and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
Visualization of the Effect on Token Length
We collect the first 50 samples from GSM8K and MATH test set in sequential order, and compare the output token length distribution between our TH2T and the original DeepSeek-R1-Distill-Qwen-7B in Fig. 9. We can see TH2T significantly reduces response length.
More Explanation on Motivation
We investigate the efficiency issue from a suspicion: Are current LRMs treating various task types equally with identical response strategy? We compare the accuracy gain and inference length scaling ratio between Short-CoT model (Qwen2.5-7B) and Long-CoT model (DeepSeek-R1-Distill-Qwen2.5-7B) on easy GSM8K and hard MATH-500 benchmarks. From Fig. 10, we obtain three findings:
Across the easy and difficult tasks, the current reasoning model shows similar token length scaling ratio (i.e., both around 400%), which verifies our initial suspicion that LRMs are allocating similar additional computing resources to tasks with varying difficulty levels.
Despite incurring minimal costs, basic LLM can still replicate the majority of correct responses (89.9%) of LRM’s on the easy tasks.
Compared with difficult tasks, the easy task only witnesses a minor accuracy gain (5.3% vs 22.4%) under the same computation resource allocation strategy, which is not what we expected. We aim to maintain this marginal performance improvement while keeping the associated token length increase at a minimal level.
These raise a reasonable speculation: LRMs fall short of perceiving task complexity, since massive inference resources on easy ones are superfluous and lead to limited performance gain. As discussed in Results on Redundancy Cognition of the main paper, we observe severe redundant structures in LRM’s responses to hard tasks, including unnecessary reflection along with a remarkably high proportion of reasoning loops in incorrect answers. These findings lead to the two objectives of this paper:
Enabling LRMs to think how to think, i.e., autonomously recognizing and adapting to variant levels of problem difficulty (results and analysis are presented in the Results of Difficulty Cognition section of the main paper).
Mitigating LRM overthinking by promoting redundancy cognition of internal reasoning structure (results and analysis are presented in the Results of Redundancy Cognition section of the main paper).
Prompts
For reproducibility, we provide the complete list of prompts adopted in our experiments.
Difficulty level evaluation
Fig. 11 shows the prompt template we used during difficulty level evaluation.
Baseline
Fig. 12 is the difficulty reminder for the baseline D-Prompt, which is also adopted in the pre-experiment. Fig. 13 is the setting of baseline NoThinking.
Design of Determinator
During Stage 2 of TH2T, we design a Determinator. To identify redundancy point, i.e., , the contribution effectiveness of reflective structures is assessed as shown in Fig. 15 and Fig. 16. Furthermore, when injecting hypnosis, the completeness requirement in Fig. 14 is adopted.
Limitations and Future Work
Our current approach primarily centers on standard fine-tuning using our generated heterogeneous dataset. While this method effectively enhances model performance within our present framework, it may not fully leverage the potential for optimized efficiency. Advanced reinforcement learning (RL)-based training strategies may offer a promising avenue for further performance improvements. However, RL can be very unstable and computationally intensive. Due to computational constraints, we defer the exploration of sophisticated RL techniques to future research.
We also recognize that theoretically, models of different series and parameter scales have distinct capability boundaries, leading to varying levels of task difficulty perception. In this phase, we introduce a common difficulty evaluation standard applicable to models at different scales. We plan to explore difficulty definition methods tailored to models of variant capability ranges in the future, aiming to advance more autonomous difficulty adaptation and perception strategies. One possible solution is to depend on the correctness ratio during sampling. However, a significant limitation of this method lies in its substantially increased computational overhead.
Case Study